# Two failure modes of deep transformers and how to avoid them: a unified theory of signal propagation at initialisation

### v2

@@ 3.2.2 Main result
3.2.3 Discussion
The phase diagram shown in Fig.˜2 exhibits a phase transition in the cosine similarity between tokens after self-attention. For small initialisation $\beta<\beta_{c}$, self-attention operates in a ‘spread attention’ phase where the attention layer effectively outputs the average of all tokens, making the pair-wise cosine similarity between tokens saturate at one, resulting in a rank-one representation matrix. In the absence of skip connections, this behaviour leads to rank collapse, which makes the transformer untrainable. This behaviour is also reminiscent of the “clustering” property analysed by [18, 27, 28].
For $\beta>\beta_{c}$, the self-attention layer preserves diversity, and hence information, among input tokens, suggesting this regime as a viable initialisation. However, the non-vanishing value of the IPR for $\beta>\beta_{c}$, Eq.˜12, reveals that self-attention is in a localised phase in this regime: it only attends to a few tokens which are determined by initialisation, rather than learnt, leading to the training instabilities observed by [11]. In other words, for $\beta>\beta_{c}$ self-attention suffers from entropy collapse, which cannot be avoided using skip connections. This behaviour is a central novelty of our analysis: [10] assumes the model operates entirely in the rank-collapse phase, thereby overlooking this distinct failure mode, while the “annealed” approximation of [13] does not capture the large-deviation behaviour underlying entropy collapse.
In a nutshell, Result˜1 implies that the only viable initialisation for self-attention is in the small-variance regime $\beta<\beta_{c}$ with skip connections to maintain information flow. To illustrate this, we train a one-layer transformer on a masked language modelling task at varying $\beta$. As shown in Fig.˜3, the two phases – spread vs. localised – exhibit distinct training behaviours. In the low-variance regime, skip connections help mitigate rank collapse, enabling successful training. In contrast, in the high-variance regime, the attention vectors’ entropy collapses, and the model fails to train effectively, as escaping the pathological initialisation and learning meaningful patterns would require prohibitively long training.
Overall, this result unifies two previously observed phenomena, namely rank and entropy collapse, within a single theoretical framework: the onset of either phenomenon is separated by a sharp phase transition, governed by the variance of the query and key weight initialisation, as parametrised by $\beta$. Before we determine the optimal strength of the skip connection in Section˜4, we analyse the backward pass through a self-attention layer.
@@ 3.3 The backward pass
Result˜2 shows that gradients can vanish under two conditions. First, we see that if $q=p$, i.e. if attention is uniform and maps all input tokens into the same output token, gradients vanish. In this case, we recover the well-known result of [10] that showed that if the inputs to the self-attention layer are already collapsed, and hence all the tokens are the same, gradients vanish. Result˜2 goes beyond their result to show that even if input tokens are diverse, i.e. $p\neq q$, gradients vanish if $\beta<\beta_{c}$ because in that regime, the inverse participation ratios $Y^{(2)}$ and $Y^{(3)}$ tend to zero as the sequence length $T\to\infty$. For long but finite sequences, finite-size effects reduce the threshold down to $\beta<\tilde{\beta}_{c}=\beta_{c}/2$ (see Section˜3.2.2). We numerically verify Result˜2 in Fig.˜3c, where we show that (1) the curves of gradient norm versus initialisation strength collapse with the scaling suggested by Result˜2, and that (2) gradients do indeed tend to zero for $\beta<\tilde{\beta}_{c}$.
Our analysis of the backward pass raises a paradox: initialise with $\beta<\beta_{c}$, and gradients vanish; initialise with $\beta>\beta_{c}$, and self-attention is stuck with entropy collapse. We discussed in the previous section that entropy collapse can only be avoided by initialising with small weights $\beta<\beta_{c}$ and adding skip connections. So where do the gradients come from? Result˜2 depends on the assumption that we are considering i.i.d. embeddings. While this assumption is true at initialisation, the embeddings do receive a non-zero gradient through the skip connections already at initialisation, independently of the value of $\beta$. The changes in the embeddings will break the i.i.d. assumption on the embeddings, and hence enable gradients to flow even in the trainable regime of $\beta<\tilde{\beta}_{c}$.
@@ 4 Full Transformer Block Analysis
We now briefly describe how to treat signal propagation through skip connections, Layer Norm, and the MLP, and then integrate our results into the simple iterative Section˜5 that practitioners can use to predict the evolution of the average cosine similarity across Transformer layers and obtain trainability diagrams that identify viable hyper-parameter choices for a given Transformer variant. We then give three applications of the algorithm in Section˜5.
Putting it all together
We derive the change to the average cosine similarity of tokens due to skip connections in Section˜B.5.1, where we find an update equation that reads
$\Phi_{\text{RES}_{\text{SA}}}(\rho)=\frac{p(\sigma_{v}^{2}+\alpha_{\text{SA}}^{2})}{\sigma_{v}^{2}\big(p+(q-p)Y^{(2)}(\beta)\big)+\alpha_{\text{SA}}^{2}q}$ (14)
where we recall that $\mathbb{E}[W_{V}^{\top}W_{V}]=\sigma_{v}^{2}I_{d}$. The update equations to describe propagation through MLP layers follows [1, 2]; for completeness, we restate these recursions in Section˜B.5.2. The activation function only enters through a kernel $f(\rho)$ that updates the overlap; for ReLU for example, we have that $f(\rho)=\pi^{-1}\left[\sqrt{1-\rho^{2}}+\rho\left(\pi-\arccos(\rho)\right)\right]$ [29]. Taking these results together, we arrive that the procedure described in Section˜5 to track the evolution of the average pairwise cosine similarity between tokens in a Transformer at initialisation. We close the paper by discussing several applications of the algorithm.
5 Applications
@@
Signal propagation in a vanilla transformer
We first used Section˜5 to analyse signal propagation in a BERT-style transformer [15]. Since BERT uses the post-norm convention for LayerNorm, we state the Section˜5 for the post-norm architecture; see Algorithm˜1 for the pre-norm version. The algorithm states the update for the average norm $q$ and average overlap $p$; the average cosine similarity can be read off after each block simply as $\rho=p/q$. Iterating the algorithm for different values of $\beta$ and $\alpha_{\text{SA}}$ finally yields the trainability diagram shown in Fig.˜1(d).
Placement of LayerNorm
[30] showed that placing LayerNorm before the self-attention layer and before the MLP greatly stabilises the training of deep transformers. Comparing signal propagation using our algorithm, we show in Fig.˜4 that rank collapse, corresponding to an average token similarity of $\langle\rho\rangle=1$, does indeed occur much later for pre-LN than for post-LN, confirming pre-LN as the more stable choice.
@@
Avoiding All Collapses: gain-controlled attention
A recent line of work has sought to alleviate rank collapse by directly modifying the self-attention layer itself. [21, 31] proposed to enforce the attention layer to be a perturbation of the identity. [21] derive an SDE description of the limiting distribution of the overlap (or neural covariance) matrix in the proportional limit where both width and depth go to infinity. To obtain this limit, they introduce a width-dependent temperature parameter and remove layer normalisation. However, layer normalisation is crucial to avoid entropy collapse, see Section˜3.2.2. Using our framework, we can show that simply removing the mean of the values along the sequence from the output of the standard self-attention layer, also explored in [31] and reminiscent of gain control in neuroscience, can be combined to great effect with either post-LN or pre-LN. In Fig.˜4, we show the theoretical prediction for the evolution of the average cosine similarity, illustrating how this modification alleviates rank collapse. Our preliminary experiments with a twenty-layer BERT-style transformer trained on TinyStories show that gain-controlled transformers succeed in regimes where vanilla attention fails, see Fig.˜C.1, encouraging further experiments at scale which are however out of the scope for the present paper.
6 Conclusions
We developed a theory for signal propagation in transformers that unifies the understanding of the two main failure modes in transformer training: rank collapse and entropy collapse. Our framework not only predicts when these phenomena occur for a given set of hyperparameters, but it also provides simple, flexible algorithms for deriving the trainability diagrams of a given architecture. Building on these insights, we find new evidence for the viability of a simple architectural modification of self-attention, the gain-controlled self-attention, that avoids both failure modes, which would be interesting to explore at scale in future work.
@@ Acknowledgements
We thank Antoine Maillard for helpful discussions. AG gratefully acknowledges funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Missione 4, Componente 1, Investimento 4.1 “Estensione del numero di dottorati di ricerca e dottorati innovativi per la Pubblica Amministrazione e il patrimonio culturale” (CUP G93C23000620003). SG gratefully acknowledges funding from the European Research Council (ERC) for the project “beyond2”, ID 101166056; from the European Union–NextGenerationEU, in the framework of the PRIN Project SELF-MADE (code 2022E3WYTY – CUP G53D23000780001), and from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research” (CUP G53C22000440006).
@@ A.1 Figure 1
Attention visualizations obtained using BertViz [32] on a single-layer, single-head Transformer at initialization. The attention maps illustrate the effect of varying $\beta$ (directly related to variance of queries/keys): in yellow, a model initialized with $\beta=0.1$ (low-variance regime, resulting in approximately uniform attention distributions), and in red, a model initialized with $\beta=1.8$ (high-variance regime, leading to sharp attention).
Trainability diagram for a 60-layer Transformer. The diagram is based on the critical value $\beta_{c}=\sqrt{2}$, which marks the threshold at which entropy collapse occurs in the first self-attention layer, assuming a sequence of orthogonal token embeddings. The residual strength threshold is defined as the smallest value of the residual scaling factor $\alpha_{\text{SA}}$ below which rank collapse (loss of representation diversity) occurs across 60 layers.
@@ Appendix B Theory Appendix
Consider the attention scores defined in Eq.˜2. By the Central Limit Theorem, they are converge in distribution to Gaussian random variables with zero mean and variance $\sigma_{a}^{2}$, where $d$ is the embedding dimension (or head dimension in the multi-head case), and $\sigma_{a}^{2}$ is determined by the initialisation of $W_{Q}$ and $W_{K}$ as described in Section˜2. Although the scores $a_{tt^{\prime}}$ are individually Gaussian, they are not independent; in fact, they are correlated. To quantify these correlations, we compute:
@@
Since the query and key weights are independently initialised with variances $\sigma_{Q}^{2}=\sigma_{K}^{2}=\nicefrac{{\sigma_{a}^{2}}}{{d}}$, this simplifies to:
$\mathbb{E}\left[a_{ts}a_{\tau\sigma}\right]=\sigma_{Q}^{2}\sigma_{K}^{2}\,(X_{t}\cdot X_{\tau})(X_{s}\cdot X_{\sigma})=\sigma_{a}^{2}q_{t\tau}q_{s\sigma}.$ (B.1)
@@ B.2.1 Computation of $Y^{(2)}(\beta)$
$\Phi_{t}(\beta,h)=\mathbb{E}\log Z_{t}(\beta,h)$ (B.2)
@@
$=\sum_{s,s^{\prime}=1}^{T}\mathbb{E}[a_{ts}a_{ts^{\prime}}]\,\mathbb{E}\left[\partial_{a_{ts^{\prime}}}\left(\frac{e^{ha_{ts}}}{\sum_{u}e^{ha_{tu}}}\right)\right]$ (B.3)
with the correlation between attention scores given by Eq.˜B.1. We assume $q_{tt}\simeq q$ for all $t$, due to concentration of measure. For the pairwise overlaps, we proceed as follows: at the first layer, all tokens are approximately orthogonal, making it safe to assume $q_{ts}\simeq 0\ \forall t\neq s\in[T]$. One then derives the update for this layer and observes that, in the limit of infinite sequence length (as we will briefly show), the update becomes independent of the indices to leading order. By repeating this argument across layers, it is therefore justified to treat the $q_{ts}$ as having a common mean $p$ together with sub-leading Gaussian fluctuations, which we neglect since our analysis focuses on the average overlap. Applying this argument, we get:
@@
$\lim_{T\to\infty}\mathbb{E}Y_{t}^{(2)}=1-\frac{1}{\sigma_{a}^{2}q\left(q-p\right)}\lim_{h\to 1}\lim_{T\to\infty}\partial_{h}\Phi_{t}(\beta,h)$ (B.4)
@@
The remaining computation amounts to evaluating a variant of the Random Energy Model (REM), but with correlated energy levels due to the structure of the $a_{ts}$ variables. This problem can be tackled using the Replica method or, alternatively, via a micro-canonical argument, as discussed in [33]. Here, we proceed with the Replica method.
@@
$=\sum_{s_{1},\dots,s_{n}=1}^{T}\exp\left(\frac{h^{2}\sigma_{a}^{2}}{2}\left(q\sum_{a,b}\mathbb{I}(s_{a}=s_{b})+qp\sum_{\begin{subarray}{c}s\neq s^{\prime}\end{subarray}}\sum_{a,b}\mathbb{I}(s=s_{a})\mathbb{I}(s^{\prime}=s_{b})\right)\right)$ (B.5)
@@
$=\sum_{Q}S(Q)\exp\left(\frac{h^{2}\sigma_{a}^{2}}{2}q\left(q-p\right)\sum_{a,b}Q_{ab}+O(n^{2})\right)$ (B.6)
@@
$\Phi_{t}(\beta,h)/N=\max\limits_{x<1}\frac{\beta^{2}h^{2}}{2}q\left(q-p\right)x+\frac{1}{x}$ (B.7)
@@
where $\beta_{c}(q,p)=\beta_{c}(h=1,q,p)=\sqrt{\frac{2}{q\left(q-p\right)}}$. We plot in Fig.˜B.1 the result of the computation and some numerical simulations.
Figure B.1: Theory and experiments ($T=10^{5}$) comparison of the computation of $Y^{(2)}(\beta)$, finite size effects are visible around the phase transition.

### v3

@@ 3.2 Forward Signal Propagation through a self-attention layer
Take-Home Message. The main implication of result˜1 is that the only viable initialisation for self-attention lies in the small-variance regime $\beta<\beta_{c}$, supplemented by skip connections to maintain information flow. To illustrate this, we train a one-layer Transformer on a masked language modelling task at varying $\beta$. As shown in fig.˜3, the two phases – spread vs. localised – exhibit distinct training behaviours. In the low-variance regime, skip connections help mitigate rank collapse, enabling successful training. In contrast, in the high-variance regime, the attention vectors’ entropy collapses, and the model fails to train effectively, as escaping the pathological initialisation and learning meaningful patterns would require prohibitively long training. Overall, this result unifies two previously observed phenomena, namely rank and entropy collapse, within a single theoretical framework: the onset of either phenomenon is separated by a sharp phase transition, governed by the variance of the query and key weight initialisation, as parametrised by $\beta$.
Sequence Geometry in Depth. We assumed i.i.d. token embeddings, which implies that before any processing the sequence forms a simple random geometry: by concentration of measure, the tokens lie near the vertices of a high-dimensional simplex, with $q_{tt}\simeq 1$ and $q_{ts}\simeq 0$. As the sequence is propagated through a Transformer at initialisation, this simplex structure is preserved in the long-sequence limit (i.e. $q_{tt}\simeq q$ and $q_{ts}\simeq p$), as shown in section˜B.2.3, although the values of $q$ and $p$ themselves evolve with depth. Thus, result˜1 provides a way to track how the geometry changes across layers: given the geometric pair $(q,p)$ at layer $l$, we plug it into the update map to obtain the new pair—and hence the updated geometry—at layer $l+1$.
Before introducing the remaining Transformer components—residual connections, MLPs, and layer normalisation—we analyse the backward pass through the self-attention layer.
@@ 3.3 The backward pass
Vanishing Gradients. Result˜2 shows that gradients can vanish under two conditions. First, we see that if $q=p$, i.e. if attention is uniform and maps all input tokens into the same output token, gradients vanish. In this case, we recover the well-known result of Noci et al. (2022) that showed that if the inputs to the self-attention layer are already collapsed, and hence all the tokens are the same, gradients vanish. Result˜2 goes beyond their result to show that even if input tokens are diverse, i.e. $p\neq q$, gradients vanish if $\beta<\beta_{c}$ because in that regime, the inverse participation ratios $Y^{(2)}$ and $Y^{(3)}$ tend to zero as the sequence length $T\to\infty$. For long but finite sequences, finite-size effects reduce the threshold down to $\beta<\tilde{\beta}_{c}=\beta_{c}/2$ (see section˜3.2). We numerically verify result˜2 in fig.˜3c, where we show that (1) the curves of gradient norm versus initialisation strength collapse with the scaling suggested by result˜2, and that (2) gradients do indeed tend to zero for $\beta<\tilde{\beta}_{c}$.
Our analysis of the backward pass raises a paradox: initialise with $\beta<\beta_{c}$, and gradients vanish; initialise with $\beta>\beta_{c}$, and self-attention is stuck with entropy collapse. We discussed in the previous section that entropy collapse can only be avoided by initialising with small weights $\beta<\beta_{c}$ and adding skip connections. So where do the gradients come from? Result˜2 relies on the simplex structure of the embeddings. While this structure holds exactly at initialisation, the embeddings receive a non-zero gradient through the skip connections already in the first backward pass, independently of the value of $\beta$. These updates break the simplex structure on which our derivations depend, and thus also invalidate the vanishing-gradient result, which applies only at initialisation.
@@ 4 Full Transformer Block Analysis
We now describe how to track signal propagation through skip connections, LayerNorm, and the MLP, and then show how these components combine into simple iterative algorithms. Such algorithms allow practitioners to predict the evolution of the average cosine similarity across Transformer layers and to generate trainability diagrams that reveal viable hyper-parameter regimes for a given Transformer architecture. An example for BERT-style post-norm Transformers is given in section˜4; additional variants, including pre-norm, are reported in section˜B.5.2.
Self-attention layer. The order parameters characterizing sequence geometry evolve in the self-attention layer as derived in section˜B.2.3:
$q_{\mathcal{S}}=\sigma_{v}^{2}\left(p+(q-p)\,Y^{(2)}(\beta,q,p)\right),\qquad p_{\mathcal{S}}=\sigma_{v}^{2}p,$ (14)
where we recall that $\mathbb{E}[W_{V}^{\top}W_{V}]=\sigma_{V}^{2}\mathbb{I}_{d}$. Taking the ratio of these expressions recovers the cosine-similarity update of result˜1.
Adding residual connections. Because $W_{V}$ is independent of the residual stream, adding a residual connection of strength $\alpha_{\text{SA}}$ contributes additively (see section˜B.5.1):
$q\leftarrow q_{\mathcal{S}}+q\,\alpha_{\text{SA}}^{2},\qquad p\leftarrow p_{\mathcal{S}}+p\,\alpha_{\text{SA}}^{2}.$ (15)
An analogous relation holds for the MLP residual path.
Adding MLPs. How the signal propagates through the MLP layers follows from the results of Poole et al. (2016); Schoenholz et al. (2016). For a two-layer ReLU MLP, the evolution of squared norms $q^{(l)}$ and pairwise inner products $p^{(l)}$ is:
$q^{(l)}=\sigma_{w}^{2}\int\mathcal{D}z\,\phi\!\left(\sqrt{q^{(l-1)}}\,z\right)^{2}+\sigma_{b}^{2},\qquad l=2,\dots,L,$
$p^{(l)}=\sigma_{w}^{2}\int\mathcal{D}\mathbf{z}\,\phi\!\left(\sqrt{q^{(l-1)}}\,z_{1}\right)\phi\!\left(\sqrt{q^{(l-1)}}\,z_{2}\right)+\sigma_{b}^{2},$
where $\phi$ is the activation function and $\mathbf{z}=(z_{1},z_{2})^{\top}$ is a pair of standard Gaussian variables with covariance $\rho^{(l-1)}=p^{(l-1)}/q^{(l-1)}$. The initial conditions after the first linear layer are $q^{(1)}=q\,\sigma_{w}^{2}+\sigma_{b}^{2}$ and $p^{(1)}=p\,\sigma_{w}^{2}+\sigma_{b}^{2}$. For a two-layer ReLU MLP, the second-layer outputs simplify to:
$q^{(2)}$ $=\tfrac{\sigma_{w}^{2}}{2}\,q^{(1)}+\sigma_{b}^{2},\quad p^{(2)}=\tfrac{\sigma_{w}^{2}}{2}\,q^{(1)}f(\rho^{(1)})+\sigma_{b}^{2},$ (16)
@@
where $f(\rho)$ captures correlations after the ReLU nonlinearity Cho and Saul (2009):
$f(\rho)=\tfrac{1}{\pi}\left(\sqrt{1-\rho^{2}}+\rho\left(\pi-\arccos(\rho)\right)\right).$
After including the MLP residual connection, the updated quantities become:
$q\leftarrow q^{(2)}+q\,\alpha_{\text{MLP}}^{2},\qquad p\leftarrow p^{(2)}+p\,\alpha_{\text{MLP}}^{2}.$ (17)
Taking their ratio gives the full block-level update of the average cosine similarity.
LayerNorm. The remaining component is layer normalisation, which simply enforces
$p\leftarrow\tfrac{p}{q},\qquad q\leftarrow 1.$
This update should be applied at the position dictated by the specific Transformer variant under consideration.
We conclude the paper by outlining several applications of the resulting algorithm.
5 Applications
@@
Signal propagation in a vanilla Transformer. We first used section˜4 to analyse signal propagation in a BERT-style Transformer (Devlin et al., 2019). Since BERT uses the post-norm convention for LayerNorm, we state the section˜4 for the post-norm architecture; see algorithm˜1 for the pre-norm version. The algorithm states the update for the average norm $q$ and average overlap $p$; the average cosine similarity can be read off after each block simply as $\rho=p/q$. Iterating the algorithm for different values of $\beta$ and $\alpha_{\text{SA}}$ finally yields the trainability diagram shown in fig.˜1(d).
Placement of LayerNorm. Xiong et al. (2020) showed that placing LayerNorm before the self-attention layer and before the MLP greatly stabilises the training of deep Transformers. Comparing signal propagation using our algorithm, we show in fig.˜4 that rank collapse, corresponding to an average token similarity of $\langle\rho\rangle=1$, does indeed occur much later for pre-LN than for post-LN, confirming pre-LN as the more stable choice.
@@
Avoiding All Collapses: gain-controlled attention. A recent line of work has sought to alleviate rank collapse by directly modifying the self-attention layer itself. Noci et al. (2023); Naderi et al. (2024) proposed to enforce the attention layer to be a perturbation of the identity. Noci et al. (2023) derive an SDE description of the limiting distribution of the overlap (or neural covariance) matrix in the proportional limit where both width and depth go to infinity. To obtain this limit, they introduce a width-dependent temperature parameter and remove layer normalisation. However, layer normalisation is crucial for preventing entropy collapse: the critical value $\beta_{c}(q,p)$ identified in result˜1 depends inversely on the squared average norm $q$. Using our framework, we can show that simply removing the mean of the values along the sequence from the output of the standard self-attention layer, also explored in (Naderi et al., 2024) and reminiscent of gain control in neuroscience, can be combined to great effect with either post-LN or pre-LN. In fig.˜4, we show the theoretical prediction for the evolution of the average cosine similarity, illustrating how this modification alleviates rank collapse. Our preliminary experiments with a thirty-layer BERT-style Transformer trained on TinyStories show that gain-controlled Transformers succeed in regimes where vanilla attention fails, see fig.˜6, encouraging further experiments at scale which are however out of the scope for the present paper. As a final note, in order to propagate the signal infinitely deep, one should simply use gain-controlled attention and initialise the MLPs at the edge of chaos.
Autoregressive Models. Our theory relies on the observation that, when starting from embeddings with a simplex structure, non-causal transformers preserve this structure in depth in the limit of infinite sequence length. In this regime, the representation dynamics can be fully described by only two parameters: the average squared norm $q$ and the average overlap $p$. A key ingredient in this result is the infinite-sequence limit itself. For autoregressive models, however, this limit introduces additional subtleties. Early tokens do not satisfy the simplex geometry because they have access to only a small fraction of the sequence context, and in this regime the geometry is more complex and two parameters are no longer sufficient. As one moves to later tokens, the geometry is progressively restored. Consequently, in causal models rank collapse does not occur at the level of the full sequence; instead, for sufficiently long sequences, it can emerge in the representations of tokens appearing later in the sequence. Our theory approximates late-position tokens, tracking how their pairwise similarity changes with depth and predicting when their representations collapse (fig.˜12).
Conclusions
We developed a theory for signal propagation in Transformers that unifies the understanding of the two main failure modes in Transformer training: rank collapse and entropy collapse. Building on these insights, we find new evidence for the viability of a simple architectural modification of self-attention, the gain-controlled self-attention, that avoids both failure modes, which would be interesting to explore at scale in future work.
@@ Acknowledgements
We thank Antoine Maillard for helpful discussions. AG gratefully acknowledges funding from Next Generation EU, in the context of the National Recovery and Resilience Plan, Missione 4, Componente 1, Investimento 4.1 “Estensione del numero di dottorati di ricerca e dottorati innovativi per la Pubblica Amministrazione e il patrimonio culturale” (CUP G93C23000620003), and also gratefully acknowledges financial support from Fondazione Zegna. SG gratefully acknowledges funding from the European Research Council (ERC) for the project “beyond2”, ID 101166056; from the European Union–NextGenerationEU, in the framework of the PRIN Project SELF-MADE (code 2022E3WYTY – CUP G53D23000780001), and from Next Generation EU, in the context of the National Recovery and Resilience Plan, Investment PE1 – Project FAIR “Future Artificial Intelligence Research” (CUP G53C22000440006).
@@ A.1 Figure 1
Attention visualizations obtained using BertViz [Vig, 2019] on a single-layer, single-head Transformer at initialization. The attention maps illustrate the effect of varying $\beta$ (directly related to variance of queries/keys): in yellow, a model initialized with $\beta=0.1$ (low-variance regime, resulting in approximately uniform attention distributions), and in red, a model initialized with $\beta=1.8$ (high-variance regime, leading to sharp attention).
(c) Trainability Diagram for a 60-Layer BERT Transformer.
The diagram is based on the critical value $\beta_{c}=\sqrt{2}$, which marks the threshold at which entropy collapse occurs in the first self-attention layer under the assumption of i.i.d. and normalised token embeddings. In post-norm architectures (such as BERT), we always have $q=1$ at the self-attention layer, so $\beta_{c}=\sqrt{2/(1-\rho)}$. Since $\rho$ increases with depth, the smallest critical value is the one at the first layer (where $\rho\simeq 0$), and we therefore use this value as a single critical threshold for the entire network (i.e. all layers are initialised with the same variance). In other Transformer variants, such as pre-norm, the critical threshold also depends on the average squared norm, so one must determine the appropriate critical value for each layer individually.
The residual-strength threshold is defined as the smallest value of the residual scaling factor $\alpha_{\text{SA}}$ below which rank collapse (loss of representation diversity) occurs across all 60 layers.
@@ Appendix B Theory Appendix
Consider the attention scores defined in eq.˜2. By the Central Limit Theorem, they converge in distribution to Gaussian random variables with zero mean and variance $\sigma_{a}^{2}$, where $d$ is the embedding dimension (or head dimension in the multi-head case), and $\sigma_{a}^{2}$ is determined by the initialisation of $W_{Q}$ and $W_{K}$ as described in section˜2. Although the scores $a_{tt^{\prime}}$ are individually Gaussian in the infinite width limit ($d\to\infty$), they are not independent; in fact, they are correlated. To quantify these correlations, we compute:
@@
Since the query and key weights are independently initialised with variances $\sigma_{Q}^{2}=\sigma_{K}^{2}=\nicefrac{{\sigma_{a}}}{{d}}$, this simplifies to:
$\mathbb{E}\left[a_{ts}a_{\tau\sigma}\right]=\sigma_{Q}^{2}\sigma_{K}^{2}\,(X_{t}\cdot X_{\tau})(X_{s}\cdot X_{\sigma})=\sigma_{a}^{2}q_{t\tau}q_{s\sigma}.$ (18)
@@ B.2.1 Computation of $Y^{(2)}(\beta)$
$\Phi_{t}(\beta,h)=\mathbb{E}\log Z_{t}(\beta,h)$ (19)
@@
$=\sum_{s,s^{\prime}=1}^{T}\mathbb{E}[a_{ts}a_{ts^{\prime}}]\,\mathbb{E}\left[\partial_{a_{ts^{\prime}}}\left(\frac{e^{ha_{ts}}}{\sum_{u}e^{ha_{tu}}}\right)\right]$ (20)
with the correlation between attention scores given by eq.˜18. We assume $q_{tt}\simeq q$ for all $t$, due to concentration of measure. For the pairwise overlaps, we proceed as follows: at the first layer, all tokens are approximately orthogonal, making it safe to assume $q_{ts}\simeq 0\ \forall t\neq s\in[T]$. One then derives the update for this layer and observes that, in the limit of infinite sequence length (as we will briefly show), the update becomes independent of the indices to leading order. By repeating this argument across layers, it is therefore justified to treat the $q_{ts}$ as having a common mean $p$ together with sub-leading Gaussian fluctuations, which we neglect since our analysis focuses on the average overlap. Applying this argument, we get:
@@
$\lim_{T\to\infty}\mathbb{E}Y_{t}^{(2)}=1-\frac{1}{\sigma_{a}^{2}q\left(q-p\right)}\lim_{h\to 1}\lim_{T\to\infty}\partial_{h}\Phi_{t}(\beta,h)$ (21)
@@
The remaining computation amounts to evaluating a variant of the Random Energy Model (REM), but with correlated energy levels due to the structure of the $a_{ts}$ variables. This problem can be tackled using the Replica method or, alternatively, via a micro-canonical argument, as discussed in Mézard and Montanari [2009]. Here, we proceed with the Replica method.
@@
$=\sum_{s_{1},\dots,s_{n}=1}^{T}\exp\left(\frac{h^{2}\sigma_{a}^{2}}{2}\left(q\sum_{a,b}\mathbb{I}(s_{a}=s_{b})+qp\sum_{\begin{subarray}{c}s\neq s^{\prime}\end{subarray}}\sum_{a,b}\mathbb{I}(s=s_{a})\mathbb{I}(s^{\prime}=s_{b})\right)\right)$ (22)
@@
$=\sum_{Q}S(Q)\exp\left(\frac{h^{2}\sigma_{a}^{2}}{2}q\left(q-p\right)\sum_{a,b}Q_{ab}+O(n^{2})\right)$ (23)
@@
$\Phi_{t}(\beta,h)/N=\max\limits_{x<1}\frac{\beta^{2}h^{2}}{2}q\left(q-p\right)x+\frac{1}{x}$ (24)
@@
where $\beta_{c}(q,p)=\beta_{c}(h=1,q,p)=\sqrt{\frac{2}{q\left(q-p\right)}}$. We plot in fig.˜5 the result of the computation and some numerical simulations.
Figure 5: Theory and experiments ($T=10^{5}$) comparison of the computation of $Y^{(2)}(\beta)$, finite size effects are visible around the phase transition.